14:20 Deployed ARGOS source and powered up the WFS

Initial bayside positions
  • 65.578, 20.482, 62.0
Hexapod positions:
UT20230413_Hexapod.png

14:37 Closed loop:
  • dist_flao2_KL_v20_atm_s1.6_L040.0_v15.0_ovfreq1700.00_sd3892.fits
  • bin1 1444 Hz, modulation 3, EMgain 600, OCAM mag 11.4
  • ACE table used: table_LBTI_ACE-AO.500m1700Hz_20230407.txt
14:58 Slope zero on OCAM. Reoptimize and reapplication of GOpt resolved.

15:01 Adjusting the bayside positions to center on SHARK camera
  • Bayside position was adjusted to 65.77, 20.283
Masterdiagnostics is producing a steam of dump files. Xianyu has already removed 1000 dump files.

15:08 Starting test by sending through the offset Z command through the software.
  • First offset 0.3mm in Z which was showed Zoffset 0.3mm in AOS and successfully moved Z bayside stage from 62.0 to 62.3
  • Sent ZOffset of -3mm, went through in AOS, moved bayside stage in steps in steps of 1mm from 62.3 to 59.3. The command on the SHARK side waited until the command completed.
15:20 Negative intensity error on OCAM
  • GOpt not working, ratio non sensical. Injection looks ok on modal plot. Tried reoptimizing but same result.
    • TN 20230413_153929
My mistake, didn't reset modulation after rebinning.
  • Target shifted -0.6mm in Y just with that incident. It appears that the absolute position was not maintained
  • While open loop +1.2 but they got the error message
  • Current bayside positions are: 65.772, 19.683, 59.3
  • Sent an 11mm ZOffset, AOS received the offset, they received the failure message, and the bayside did not move as expected
15:40 Moving back to nominal Z
  • They sent a +2.7 Z offset command which worked nominally
15:44 Resending presetAO in advance of their optimization. GOpt working this time
  • PSF is where it was before, no shift. Bayside positions: 65.772, 19.683, 62.001
  • They are optimizing camera
16:09 Starting template test

16:33 Zero Intensity on OCAM
  • TN 163359
  • Rebinning and Reoptimizing. Looked resolved but repotimize had another zero intensity (14:36) Zero intensity again during the ho2 second pass optimization after recovery. Removed some thin shell offload to get the source back on the hotspot
  • Rebinned and reoptimized and we are back in business.

-- JenniferPower - 13 Apr 2023 *
Topic revision: r2 - 13 Apr 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback